OpenStudy (anonymous):

can someone teach me how to figure out when one equation system doesn't have a solution.

OpenStudy (saifoo.khan):

please elaborate the question!

OpenStudy (anonymous):

in other words, which of the following equation systems cannot be solved

OpenStudy (anonymous):

when the slope of the equations are the same but the y-intercepts are different

OpenStudy (saifoo.khan):

with same slopes.. and different y intercpts.

OpenStudy (anonymous):

dont get it. With the equations above. Can you teach me

OpenStudy (saifoo.khan):

sure i can.

OpenStudy (saifoo.khan):

in the picture all can be solved.

OpenStudy (anonymous):

the last one can't be solved.

OpenStudy (saifoo.khan):

Lol, yea. i didnt saw that.

OpenStudy (saifoo.khan):

sorry!

OpenStudy (anonymous):

Its ok dude Acoording to what I understand. I should put the equations into the form y=mx+b and compare slopes and y intencepts, right?

OpenStudy (saifoo.khan):

look: x + 3y = 6 3x+9y = 7 if u look @ the second equation, thats the triple of the first. ================================================== Oh Yea u got the point right!

OpenStudy (anonymous):

let me try; if I get stacked, Ill ask you for help. Thanks

OpenStudy (saifoo.khan):

Sure! :]

OpenStudy (anonymous):

I got it, thank you all :)

OpenStudy (saifoo.khan):

Uw. :]